OpenStudy (anonymous):

GEOMETRIC SEQUENCE HELP: What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250? So far I've got to find r: a8=10r^(8-1) an=10r^7 781250=10r^7

OpenStudy (anonymous):

any help is appreciated

OpenStudy (wolf1728):

log 781250 = 7 * log 10r 7 = log 781250 / log 10r 7 = 5.8927900304 / log 10r log 10r = 5.8927900304 / 7 log 10r = 0.8418271472 10r = 6.947477472 r = .6947477472

OpenStudy (anonymous):

erm although that may be right I thought I was suppose to get r=5 ? without the use of logs

OpenStudy (wolf1728):

If r = 5 then 10r ^7 ?= 781250 10r = 50 50^7 = 781,250,000,000 5^7 = 78,125 I think that 10r is creating problems

OpenStudy (anonymous):

okay well im using this an=a1*r^(n-1) that's were i got everything from im just having trouble finding the final answer of r

OpenStudy (wolf1728):

what does an mean and also a1?

OpenStudy (anonymous):

an number of terms in the sequence and a1 is the first term

OpenStudy (wolf1728):

Okay how about explaining how you got this? 781250=10r^7 And where does that 10 come from?

OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (anonymous):

10=a1 8=n 781250=an

OpenStudy (anonymous):

a₈ = a₁*r^(8 - 1) ==> 781250 = 10r^7 r = (781250/10)^(1/7) = 5. that's the answer apparently but this is the part I'm not understanding, this isn't my work

OpenStudy (wolf1728):

Wonder if the formula written like this? 781,250=(10*r)^7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):